OpenStudy (anonymous):

add and simplify: ((-1)/x)+((2/(x^2+1))+((1/(x^3+x))

OpenStudy (anonymous):

sorry it's actually ((-1)/x)+2/(x^2+1)+1/(x^3+x)

OpenStudy (anonymous):

is this what you mean? \[-\frac{ 1 }{ x } +\frac{ \frac{ 2 }{ x^{2}+ 1 }+1 }{ x^{3}+x }\]